OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (3, 4) and (–3, –8) in point-slope form. Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (amistre64):

subtract your points; and stack y over x to get the slope ( 3, 4 ) -(-3,-8) --------- 6 ,12 ; slope = y/x = 12/6 = 2 y-Py = m(x-Px) ; P(3,4) y-4 = 2(x-3)